# FREA: Feasibility-Guided Generation of Safety-Critical Scenarios with Reasonable Adversariality

### v2

@@ 1 Introduction
Rapid advancements in AI-driven autonomous driving technologies have enabled autonomous vehicles (AVs) to operate efficiently, safely, and comfortably in most daily scenarios [1]. However, their performance in the complex real world remains uncertain due to safety-critical yet rare scenarios, which are hard to collect at scale. Accurately generating these safety-critical scenarios and evaluating the AV performance is crucial for advancing autonomous driving systems [2].
@@ 3.2 Feasibility-Dependent Adversarial Objective of CBV
As the BVs follow deterministic policy, the optimal feasible action-value function of AV can be achieved by AV’s current state and next state, while the action of the CBV and AV mainly influence state transition (see Section A.2 for proof).
@@
1:Offline Part (Section 3.1)
2:Initialize feasibility value networks $V_{h}$, $Q_{h}$.
3:# Optimal feasible value function learning
4:for each gradient step do
5:     Update $V_{h}$ using Eq. 3
6:     Update $Q_{h}$ using Eq. 4
7:end for
8:Online Part (Section 3.2)
9:Initialize policy parameters $\theta_{0}$, reward value function parameters $\psi_{0}$
10:for $k=0,1,2,\dots$ do
11:     Collect set of trajectories $\mathcal{B}_{k}=\{\tau_{i}\}$ with policy $\pi_{\theta_{k}}$, where $\tau_{i}$ is a $T$-step episode.
12:     # Advantage calculating
13:     Compute reward advantage $A_{r}^{\pi_{\theta_{k}}}(s,a)$, using generalized advantage estimator (GAE [23]).
14:     Compute feasibility advantage using Eq. 9.
15:     Derive overall advantage using Eq. 6
16:     # Value function learning
17:     Fit reward value function, by Smooth L1 Loss.
18:     # Policy learning
19:     Update the policy parameters $\theta$ by maximizing Eq. 5.
20:end for
@@ 4.2 Largest Feasible Region Learning of AV
According to the loss functions for the feasible value functions detailed in Eqs. 3 and 4, learning incorporates the constraint violation function $h(s^{\textnormal{AV}})$. A straightforward definition of $h(s^{\textnormal{AV}})$ is $\min_{i\in\{1,\ldots,N\}}d(\textnormal{AV},\textnormal{BV}^{i})$, where $d(\textnormal{AV},\textnormal{BV}^{i})$ represents the minimum distance between the bounding boxes of AV and ${\textnormal{BV}}^{i}$. However, since the minimum distance between boxes is non-negative, we adopt a sparse design for $h(s^{\textnormal{AV}})$, which helps to maintain a stable training balance of positive and negative samples, as advised by [15]:
@@
Figure 2: Visualization of well-trained LFR $\left\{s^{\textnormal{AV}}|V^{\ast}_{h}(s^{\textnormal{AV}})\leq 0\right\}$ in various traffic scenarios.
@@ 4.3 Quantitative Analysis of Generated Scenarios
To verify that FREA can generate adversarial yet AV feasible scenarios compared to baseline methods, we quantified the near-miss events and the AV’s feasibility across various CBV methods. To ensure consistency and fairness in our quantitative analysis, we employed the same surrogate AV method, Expert [26], throughout the experiments. Further generalization analysis about the surrogate AV methods is discussed in Section 4.4.
@@
As shown in Table 1 and Figure 4, without the feasibility constraint, the PPO and KING [5] lead to many severe collisions, indicative of excessive adversarial behavior. In contrast, under the feasibility constraint, both FPPO-RS and FREA significantly reduce the severity of these events. However, while addressing AV’s feasibility, FPPO-RS struggles with hard constraints, which leads to many infeasible events. Conversely, FREA adeptly balances adversarial intensity with AV’s feasibility, thereby achieving the least severe collisions.
@@
Town05 intersections Town02 intersections
CBV Feasibility CR (↓) IR (↓) ID (↓) CR (↓) IR (↓) ID (↓)
@@
Town05 intersections Town02 intersections
CBV Surr. AV AV CR (↓) OR (↓) RF (↓) UC (↓) TS (↓) OS (↑) CR (↓) OR (↓) RF (↓) UC (↓) TS (↓) OS (↑)
@@ 4.5 Analysis of Representative Scenarios
To confirm the validity of scenarios generated by FREA, we analyze representative cases. Figure 5(a) illustrates a safety-critical scenario at an intersection where the AV, attempting a right turn, is cut off by a faster CBV also turning right prematurely, forcing the AV to stop and wait. Subsequently, another vehicle to the right of the AV is selected as the next CBV for further attack. Figure 5(b) depicts a scenario on a straight road where multiple CBVs sequentially adopt different strategies to attack the AV. These examples demonstrate that FREA is capable of creating a series of continuous, adversarial, yet AV-feasible scenarios in complex traffic environments. (Further analysis can be found in Appendix F)
(a) Intersection attacks
(b) Straight road attacks
@@ 5 Conclusion
In conclusion, FREA employs the concept of the Largest Feasible Region (LFR) of the AV to establish the upper bound of adversariality in safety-critical scenario generation. This approach aims to maximize adversariality while ensuring AV’s feasibility. Specifically, FREA employs a two-stage framework that begins by pre-training the AV’s optimal feasible value network using offline datasets to establish the LFR. Subsequently, it learns a reasonable adversarial policy for Critical Background Vehicles (CBVs) based on a feasibility-dependent adversarial objective, thereby creating adversarial yet AV-feasible safety-critical scenarios. Experimental results illustrate that the scenarios generated by FREA contain considerable near-miss events while effectively reducing unavoidable collision events. Moreover, guided by the AV’s feasibility and the goal-based adversarial reward, FREA demonstrates enhanced robustness and generalization ability in AV testing across various environments and surrogate AV methods.
@@ Appendix D Robustness of AV Pre-trained with Various CBV Methods
Town05 intersections Town02 intersections
Surr. CBV CBV CR (↓) OR (↓) RF (↓) UC (↓) TS (↓) OS (↑) CR (↓) OR (↓) RF (↓) UC (↓) TS (↓) OS (↑)

### v3

@@ 1 Introduction
Rapid advancements in AI-driven autonomous driving technologies have enabled autonomous vehicles (AVs) to operate efficiently, safely, and comfortably in most daily scenarios [1]. However, their performance in the complex real-world remains uncertain due to safety-critical yet rare scenarios, which are hard to collect at scale. Accurately generating these safety-critical scenarios and evaluating the AV performance is crucial for advancing autonomous driving systems [2].
@@ 3.2 Feasibility-Dependent Adversarial Objective of CBV
1:Offline Part (Section 3.1)
2:Initialize feasibility value networks $V_{h}$, $Q_{h}$.
3:for each gradient step do
4:     Update $V_{h}$ using Eq. 3# Optimal feasible state-value function learning
5:     Update $Q_{h}$ using Eq. 4# Optimal feasible action-value function learning
6:end for
7:Online Part (Section 3.2)
8:Initialize policy parameters $\theta_{0}$, reward value function parameters $\psi_{0}$
9:for $k=0,1,2,\dots$ do
10:     Collect set of trajectories $\mathcal{B}_{k}=\{\tau_{i}\}$ with policy $\pi_{\theta_{k}}$, where $\tau_{i}$ is a $T$-step episode.
11:     Compute reward advantage $A_{r}^{\pi_{\theta_{k}}}(s,a)$, using generalized advantage estimator (GAE [23]).
12:     Compute feasibility advantage using Eq. 9.
13:     Derive overall advantage using Eq. 6# Advantage calculating
14:     Fit reward value function, by Smooth L1 Loss.  # Value function learning
15:     Update the policy parameters $\theta$ by maximizing Eq. 5.  # Policy learning
16:end for
@@
As the BVs follow deterministic policy, the optimal feasible action-value function of AV can be achieved by AV’s current state and next state (see Section A.2 for proof).
@@ 4.2 Largest Feasible Region Learning of AV
According to Eqs. 3 and 4, learning the optimal feasible value functions incorporates the constraint violation function $h(s^{\textnormal{AV}})$. A simple definition of $h(s^{\textnormal{AV}})$ is $\min_{i\in\{1,\ldots,N\}}d(\textnormal{AV},\textnormal{BV}^{i})$, where $d(\textnormal{AV},\textnormal{BV}^{i})$ represents the minimum distance between the bounding boxes of AV and ${\textnormal{BV}}^{i}$. However, since the minimum distance is non-negative, we adopt a sparse design for $h(s^{\textnormal{AV}})$, which helps to maintain a stable training balance of positive and negative samples, as advised by [15]:
@@
Figure 2: Visualization of well-trained LFR in various scenarios.
@@ 4.3 Quantitative Analysis of Generated Scenarios
To verify that FREA can generate adversarial yet AV feasible scenarios, we quantified the near-miss events and the AV’s feasibility across various CBV methods. To ensure consistency and fairness in our quantitative analysis, we employed the same surrogate AV method, “Expert” [26], throughout the experiments. Further analysis about the surrogate AV methods is discussed in Section 4.4.
@@
CBV Feasibility Town05 intersections Town02 intersections
CR (↓) IR (↓) ID (↓) CR (↓) IR (↓) ID (↓)
@@
As shown in Figure 4 and Table 1, without the feasibility constraint, the PPO and KING [5] lead to many severe collisions, indicative of excessive adversarial behavior. In contrast, under the feasibility constraint, both FPPO-RS and FREA significantly reduce the severity of these events. However, FPPO-RS still struggles with hard constraints, leading to many infeasible events. Conversely, FREA balances adversarial intensity with AV’s feasibility, thereby achieving the least severe collisions.
@@
CBV Surr. AV AV Town05 intersections Town02 intersections
CR (↓) OR (↓) RF (↓) UC (↓) TS (↓) OS (↑) CR (↓) OR (↓) RF (↓) UC (↓) TS (↓) OS (↑)
@@ 4.5 Analysis of Representative Scenarios
To confirm the validity of scenarios generated by FREA, we analyze representative cases. Figure 5(a) illustrates a scenario at an intersection where the AV, attempting a right turn, is cut off by a faster CBV also turning right prematurely, forcing the AV to stop and wait. Subsequently, another vehicle to the right of the AV is selected as the next CBV for further attack. Figure 5(b) depicts a scenario on a straight road where multiple CBVs sequentially adopt different strategies to attack the AV. These examples demonstrate that FREA is capable of creating a series of continuous, adversarial yet AV-feasible scenarios in complex traffic environments. (Further analysis can be found in Appendix F)
@@ 5 Conclusion
In conclusion, FREA employs the concept of the Largest Feasible Region (LFR) to establish the upper bound of adversariality in safety-critical scenario generation. This approach aims to maximize adversariality while ensuring AV’s feasibility. Specifically, FREA employs a two-stage framework that begins by pre-training the AV’s optimal feasible value network using offline datasets to establish the LFR. Then, it learns a reasonable adversarial policy for Critical Background Vehicles (CBVs) based on a feasibility-dependent adversarial objective, thereby creating adversarial yet AV-feasible scenarios. Experimental results illustrate that the scenarios generated by FREA contain considerable near-miss events while effectively reducing unavoidable collision events. Moreover, guided by the AV’s feasibility and the goal-based adversarial reward, FREA demonstrates enhanced robustness and generalization ability in AV testing across various environments and surrogate AV methods.
@@ Appendix D Robustness of AV Pre-trained with Various CBV Methods
Surr. CBV CBV Town05 intersections Town02 intersections
CR (↓) OR (↓) RF (↓) UC (↓) TS (↓) OS (↑) CR (↓) OR (↓) RF (↓) UC (↓) TS (↓) OS (↑)